Consider a generic element A. It exists in the gas phase as a triatomic molecule A3 or as a diatomic molecule A2.

Calculate ΔSo in J/K for the reaction: 2A3 → 3A2 using the following values

So = 207.9 J/(mol.K) for A2(g) and 236.5 J/(mol.K) for A3(g) at 25 oC.
Answer #1

given reaction is 2A3  \rightarrow 3A2

media%2F92e%2F92ec6800-9836-4fee-9d3a-16 S0 A3= 236.5J/K

media%2Faa8%2Faa8b7ed9-9ba2-4420-ae9c-54 S0 A2= 207.9 J/K

http://latex.codecogs.com/gif.latex?media%2F0ad%2F0ad624e9-50dd-4e34-bdca-39 S0 = media%2F576%2F576ff83e-529f-42bc-a3db-d9media%2F555%2F555b5eb8-93de-4087-b3d2-26 S0 (Product) - media%2Fa4f%2Fa4f727d6-76da-42ca-8478-97media%2Fc07%2Fc0748e38-27d6-48b8-96e4-8e S0 (Reactant)

media%2F329%2F32963b05-f867-473d-a106-71 S0 = [(3media%2F442%2F442d7089-c38d-45d1-8b70-24 S0 A2)] - [(2Xmedia%2Ff7f%2Ff7f18fc0-6652-49b0-8a66-b5 S0 A3)] substitute above values in this formula

= [  (3media%2F4b6%2F4b60efef-83c2-4bc7-8a60-3a207.9 )] - [(2x236.5)]

= [623.7] - [473] = 150.7

ans media%2Fe70%2Fe709f9f6-e6b5-4c05-bbc6-79 S0 = 150.7 J/K for given reaction